I have a algebra problem that, I need help solving. Can you please help me.
sqrt (x+20) -x=0

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

adding x ___ sqrt(x+20) = x

squaring ___ x + 20 = x^2

subtracting (x + 20) ___ 0 = x^2 - x - 20

factoring ___ 0 = (x - 5)(x + 4)

Answer by stanbon(75887)

sqrt (x+20) -x=0
----
sqrt(x+20) = x
----
Square both sides:
x+20 = x^2
----
Rearrange:
x^2 -x - 20 = 0
---------
Factor:
(x-5)(x+4) = 0
----
x = 5 or x = -4
-----
Check:
Can x = 5 ?
sqrt(5+20)-5 = 0
5 - 5 = 0
------------------------
Can x = -4 ?
sqrt(-4+20)-(-4) = 0
sqrt(16) + 4 = 0
4 + 4 = 0
False
=======
Only solution : x = 5
========================
Cheers,
Stan H.
=======================